OpenStudy (anonymous):

Use the Pythagorean theorem to find the length of the hypotenuse with one side measuring 16 feet and the other measuring 7 feet. 21.34 feet 17.46 feet 14.39 feet 16.78 feet

OpenStudy (anonymous):

a^2=b^2+c^2 a^2=16^2+7^2